40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 13 00:11:43 2009, Jenne, Update, IOO, Mode Cleaner Unlock 
    Reply  Thu Aug 13 09:22:06 2009, rana, Update, PEM, ranger 
Message ID: 1897     Entry time: Thu Aug 13 09:22:06 2009     In reply to: 1895
Author: rana 
Type: Update 
Category: PEM 
Subject: ranger 

Rana, Jan, Jenne

We noticed that the Ranger data was all bogus at low frequencies. So we checked it and found that the proper procedure had not been used when changing it from horizontal to vertical last week. So the huddle test data from the weekend is not valid for the ranger; we will have to repeat it sometime.

So we used the manual, and extended the hanger rod on top of the Ranger to free the mass. It now has good response and coherence with the Guralps down to 0.1 Hz. See attached plot soon.

 

ELOG V3.1.3-